determine the amount of change in coulomb in 1 gm ion of N3-

Grade:12

1 Answers

Arun
6 years ago

Q=ne

n- represents no of electron.

e -represents the charge of electron.

1 mole - 14 gm of N3- ions

1 gm - 1/14 mole

1 gm- 1/14 * 6.022*10^23 atoms

1 atom of N3- have (7 +3) =10 electrons

therefore, 1 gm n^3- have 10* 1/14*6.022*10^23 electron

Q = (10*1/14*6.022*10^23)*(1.6*10^-19)= Answer